The U.S. Department of Agriculture (USDA) is reviewing a proposal to pay farmers a price support of $3.00 per bushel for a given commodity. The market demand for the commodity is given by the equation QD = 16 – 4P, where Q is in billions of bushels. The market supply of the commodity is given by the equation QS = −2 + 2P. Rather than buying surplus production, the USDA will require that farmers only produce the quantity demanded at the support price. In other words, the USDA will pay farmers not to produce any surplus production at the support price. Assuming that the proposal is enacted, the total cost to the taxpayer is estimated to be __________ billion:

  • $9.75

  • $10.50

  • $12.75

  • $0.00

  • $12.00
Answer #1

Answer:

The total cost to the taxpayer is estimated to be $12 billion because market price of the commodity is P=$3 and market demand Q=4 which is obtained from QD=QS=>16-4P=-2+2P=>P=3 and Q=4.

Thus if the USDA will pay farmers not to produce any surplus production at the support price then USDA will buy only 4 bushel of the commodity at P=$3.

Therefore, total cost of the purchase=P*Q=$3*4=$12.